A 3.00 kg rock is dropped from the top of a 12.0 m high cliff. Using the conservation of energy, find the following. 


d. The gravitational potential energy after it has fallen 6.00 m. 

e. The kinetic energy after it has fallen 6.00 m. 

f. The gravitational potential energy just as it hits the ground. 



1
Expert's answer
2022-04-24T17:18:50-0400

d.

"U=mg(h-h_1)=180~J."

e.

"E=\frac m2\sqrt{2g(h-h_1)}^2=180~J,"

f.

"U=mg(h-h)=0~J."